OpenStudy (anonymous):

Use the distributive property to help you solve the equation 9(w-7)=18

OpenStudy (anonymous):

9w-63=18, 9w=18+63, 9w=81, w=9

OpenStudy (anonymous):

9w-63=18 ===> 9w=81 ===> w=9